OpenStudy (anonymous):

Evaluate the expression -√0.04 0.02 0.2 -0.02 -0.2 (I chose D is that right?)

OpenStudy (anonymous):

b

OpenStudy (anonymous):

D is correct.

OpenStudy (anonymous):

Are you sure you guys just gave me two different answers

OpenStudy (anonymous):

\[-\sqrt{0.04}=-\sqrt{4*10^{-2}}=-2*10^{-1}=-0.2\]

OpenStudy (anonymous):

I'm not always right, but I'm pretty sure this one is....

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Thank you!